# Haoran Guan

### v1

@@
Abstract
In this work, we focus on improving LU-CholeskyQR2 LUChol . Compared to other deterministic and randomized CholeskyQR-type algorithms, it does not require a sufficient condition on $\kappa_{2}(X)$ for the input tall-skinny matrix $X$, which ensures the algorithm’s safety in most real-world applications. However, the Cholesky factorization step may break down when the $L$-factor after the LU factorization of $X$ is ill-conditioned. To address this, we construct a new algorithm, LU-Householder CholeskyQR2 (LHC2), which uses HouseholderQR to generate the upper-triangular factor, thereby avoiding numerical breakdown. Moreover, we utilize the latest sketching techniques to develop randomized versions of LHC: SLHC and SSLHC. We provide a rounding error analysis for these new algorithms. Numerical experiments demonstrate that our three new algorithms have better applicability and can handle a wider range of matrices compared to LU-CholeskyQR2. With the sketching technique, our randomized algorithms, SLHC2 and SSLHC3, show significant acceleration over LHC2. Additionally, SSLHC3, which employs multi-sketching, is more efficient than SLHC2 and exhibits better numerical stability. It is also robust as a randomized algorithm.
Keywords:
QR factorization Rounding error analysis LU-CholeskyQR Numerical linear algebra
MSC:
65F30 15A23 65F25 65G50
1 Introduction
QR factorization is one of the most important components of numerical linear algebra and is widely used in both academia and industry (see 2018robust ; halko2011 ; 2010 ; 2020 ; 1989 for more details). Among the various algorithms for QR factorization, CholeskyQR has gained popularity in recent years due to its ability to balance efficiency and accuracy. CholeskyQR exclusively utilizes BLAS3 operations and requires only simple reductions in parallel environments, which is a significant advantage compared to other algorithms such as HouseholderQR, CGS (MGS), and TSQR ballard2011 ; 2011 ; MatrixC ; Higham ; Communication ; Numerical .
1.0.1 Deterministic CholeskyQR-type algorithms
CholeskyQR primarily focuses on the QR factorization of tall-skinny matrices. $X\in\mathbb{R}^{m\times n}$ and $m\geq n$, the Gram matrix $B$ is computed first, and the upper-triangular $R$-factor is obtained through Cholesky factorization. The orthogonal factor $Q\in\mathbb{R}^{m\times n}$ can then be derived using $Q=XR^{-1}$. CholeskyQR is outlined in Algorithm 1. Due to the requirement for Cholesky factorization in the algorithm, the input $X$ must satisfy $rank(X)=n$, which is a sufficient condition for all CholeskyQR-type algorithms.
@@ 1.0.1 Deterministic CholeskyQR-type algorithms
In practical applications, CholeskyQR is seldom used alone due to its lack of orthogonality. To enhance numerical stability in both orthogonality and residuals, CholeskyQR2 2014 ; error was developed by repeating CholeskyQR twice. This algorithm is outlined in Algorithm 2. To address the issue of ill-conditioned matrices, Shifted CholeskyQR (SCholeskyQR) Shifted introduces a shifted term $s$ during the Cholesky factorization step to avoid numerical breakdown, as detailed in Algorithm 3. Researchers have combined CholeskyQR2 with Shifted CholeskyQR to create Shifted CholeskyQR3 (SCholeskyQR3), which ensures orthogonality, as shown in Algorithm 4. In our recent work Columns , we define a new $g$-norm of $X$ and improve the original shifted term associated with $\|X\|_{2}$ by using a smaller term based on $\|X\|_{g}$. This modification expands the computational range of $\kappa_{2}(X)$ for Shifted CholeskyQR3. With the definition of the $g$-norm, we also analyze the case when $X$ is sparse CSparse . Further analysis of deterministic CholeskyQR-type algorithms can be found in Prob ; 2016 ; Mixed .
@@
1.0.2 LU-CholeskyQR2 and its bottleneck
Among all the deterministic CholeskyQR-type algorithms, LU-CholeskyQR LUChol is particularly noteworthy. It has the unique advantage of not imposing a restriction on $\kappa_{2}(X)$, which is crucial for real-world applications in industry. LU-CholeskyQR combines LU factorization with CholeskyQR, as shown in Algorithm 5. Here, $L\in\mathbb{R}^{m\times n}$ is a unit tall-skinny lower triangular matrix, and $U\in\mathbb{R}^{n\times n}$ is an upper triangular matrix when $X\in\mathbb{R}^{m\times n}$ is tall-skinny. A CholeskyQR step is performed afterwards to form LU-CholeskyQR2, ensuring orthogonality, as detailed in Algorithm 6.
@@ 1.0.2 LU-CholeskyQR2 and its bottleneck
In practice, the $L$-factor after LU factorization is often well-conditioned, even when $X$ is ill-conditioned. However, there are many real-world scenarios that can lead to an ill-conditioned $L$, such as certain lower-triangular matrices and diagonally dominant matrices with many negative elements. These situations frequently arise in linear equations, information theory, and systems theory. The primary bottleneck of LU-CholeskyQR2 is the potential for numerical breakdown in these cases, which is a significant problem that remains to be addressed or improved.
1.0.3 Randomized algorithms and our contributions
In recent years, randomized numerical linear algebra halko2011 ; 2020 has become a hot topic worldwide. A new randomized technique called sketching Ach has been widely applied to various problems. The sketching technique primarily aims to reduce computational costs by replacing the original large matrices with alternative matrices of smaller sizes after some preconditioning steps. The sketching matrix is typically chosen to be an $\epsilon$-subspace embedding or a linear map to a lower-dimensional space, preserving the inner products and norms of all vectors within the subspace up to a factor of $\sqrt{1+\epsilon}$, where $10\leq\epsilon<$ 1. Several existing methods for sketching matrices include Gaussian Sketch, SRHT, and Count Sketch rgs ; pmlr ; Fast ; Estimating ; pylspack . Research on randomized CholeskyQR has also emerged. Y. Fan and his collaborators Novel proposed the initial version of randomized CholeskyQR, while O. Balabanov Randomized provided a detailed analysis of CholeskyQR-type algorithms with various randomized strategies, including sketching. A recent work by A. J. Higgins and his collaborators Householder introduced a novel method called multi-sketching, which employs two different sketching steps consecutively. This work developed a new algorithm called randomized Householder QR, utilizing HouseholderQR to replace the original generation step of the $R$-factor in CholeskyQR to avoid numerical breakdown. The multi-sketching technique is designed to accelerate the entire algorithm. The corresponding algorithm is outlined in Algorithm 7. Here, $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ and $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ are sketching matrices, with $n\leq s_{2}\leq s_{1}\leq 1$. To ensure good orthogonality, a CholeskyQR step is added afterwards, resulting in the Rand.Householder-Cholesky algorithm, as shown in Algorithm 8. They provide a clear rounding error analysis of the algorithms with two $\epsilon$-subspace embeddings, and numerical experiments demonstrate that their randomized Householder-CholeskyQR is more efficient and numerically stable than the algorithms in Randomized .
@@ 1.0.3 Randomized algorithms and our contributions
1.0.4 Outline of this work
Our work is organized as follows: We present existing results from previous studies in Section 2. In Sections 3 and 4, we introduce our improved deterministic and randomized algorithms, along with their corresponding rounding error analyses regarding orthogonality and residuals. The primary numerical experiments and comparisons are detailed in Section 5. Finally, we conclude with a summary of the results in Section 6.
2 Literature review
@@
2.1 Rounding error analysis
@@
Lemma 1 (Weyl’s Theorem MatrixC )
@@
Lemma 2 (Rounding error in matrix multiplication Higham )
For $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times p}$, the error in computing the matrix product $C=AB$ in floating-point arithmetic is bounded by
@@
Lemma 3 (Rounding error in HouseholderQRHigham )
@@
To learn more about the rounding error analysis of matrix perturbation, readers can refer to Improved ; Perturbation ; Backward for more details.
2.2 Matrix sketching
@@
2.2.1 Single-sketching
In this section, we introduce the concept of sketching with the definition of an $\epsilon$-subspace embedding rgs ; Fast ; 4031351 .
Definition 1 ($\epsilon$-subspace embedding)
@@ 2.2.1 Single-sketching
$|\langle x,y\rangle-\langle\Omega x,\Omega y\rangle|\leq\epsilon\|x\|_{2}\|y\|_{2}.$ (1)
@@
In the following, we list some properties based on Definition 1.
Lemma 4 (Single-sketching of the vector)
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, then for any $x\in\mathcal{K}$, we can have
$\sqrt{1-\epsilon}\cdot\|x\|_{2}\leq\|\Omega x\|_{2}\leq\sqrt{1+\epsilon}\cdot\|x\|_{2}.$ (2)
Lemma 4 focuses on the sketching of vectors. Below, we list the properties of the matrix after sketching.
Lemma 5 (Single-sketching of the matrix)
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, if $T$ is a matrix whose columns form a basis of $\mathcal{K}$, we can have
$\sqrt{1-\epsilon}\cdot\|T\|_{2}\leq\|\Omega T\|_{2}\leq\sqrt{1+\epsilon}\cdot\|\Omega T\|_{2},$ (3)
$\sqrt{1-\epsilon}\cdot\|T\|_{F}\leq\|\Omega T\|_{F}\leq\sqrt{1+\epsilon}\cdot\|\Omega T\|_{F},$ (4)
$\sqrt{1-\epsilon}\cdot\sigma_{min}(T)\leq\sigma_{min}(\Omega T)\leq\|\Omega T\|_{2}\leq\sqrt{1+\epsilon}\cdot\|T\|_{2}.$ (5)
@@
In Lemma 5, equations (3) and (5) play a crucial role in our analysis. Equation (4) can be easily derived from (3). Below, we list some existing results related to condition numbers.
Lemma 6 (Condition number of the single-sketching)
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-subspace embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, if $T$ is a matrix whose columns form a basis of $\mathcal{K}$, we can have
$\frac{\sigma_{min}(\Omega T)}{\sqrt{1+\epsilon}}\leq\sigma_{min}(T)\leq\|T\|_{2}\leq\frac{\|\Omega T\|_{2}}{\sqrt{1-\epsilon}}.$ (6)
That is,
$\kappa_{2}(T)\leq\sqrt{\frac{1-\epsilon}{1+\epsilon}}\cdot\kappa_{2}(\Omega T).$ (8)
@@
The results mentioned above pertain to $\epsilon$-subspace embeddings, which require knowledge of the subspace $\mathcal{K}$. In contrast, a concept known as the $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding rgs does not require this knowledge.
Definition 2 ($(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding)
@@
Typical $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embeddings include both dense and sparse cases, which correspond to different sketching methods. A sketching matrix is defined as $\Omega=\frac{1}{s}G$, where $G\in\mathbb{R}^{s\times m}$ is a Gaussian matrix. The sketching size $s$ 4031351 satisfies
$s=\eta\cdot\frac{logn\cdot log{\frac{1}{p}}}{\epsilon^{2}}.$ (9)
@@
Regarding $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embeddings, researchers often use CountSketch Low . If $\Omega\in\mathbb{R}^{s\times m}$ is a sketching matrix for CountSketch, then, according to frequent ; tool , $s$ needs to satisfy
$s\geq\frac{n^{2}+n}{\epsilon^{2}p}.$ (10)
2.2.2 Multi-sketching
In Householder , a technique called multi-sketching is introduced, which involves applying two steps of sketching consecutively. Below, we list some properties of the multi-sketching technique.
Lemma 7 (Multi-sketching of the vector)
If $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ is a $(\epsilon_{2},p_{2},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$ and they are generated independently, then for any $n$-dimensional subspace $\mathcal{K}\subset\mathbb{R}^{m}$ and any $x\in\mathcal{K}$, we can have
$\sqrt{1-\epsilon_{s}}\cdot\|x\|_{2}\leq\|\Omega_{2}\Omega_{1}x\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|x\|_{2}$ (11)

### v2

# Deterministic and randomized LU-Householder CholeskyQR

@@
Abstract.
In this work, we focus on improving LU-CholeskyQR2 [34]. Compared to other deterministic and randomized CholeskyQR-type algorithms, it does not require a sufficient condition on $\kappa_{2}(X)$ for the input tall-skinny matrix $X$, which ensures the algorithm’s safety in most real-world applications. However, the Cholesky factorization step may break down when the $L$-factor after the LU factorization of $X$ is ill-conditioned. To address this, we construct a new algorithm, LU-Householder CholeskyQR2 (LHC2), which uses HouseholderQR to generate the upper-triangular factor, thereby avoiding numerical breakdown. Moreover, we utilize the latest sketching techniques to develop randomized versions of LHC: SLHC and SSLHC. We provide a rounding error analysis for these new algorithms. Numerical experiments demonstrate that our three new algorithms have better applicability and can handle a wider range of matrices compared to LU-CholeskyQR2. With the sketching technique, our randomized algorithms, SLHC2 and SSLHC3, show significant acceleration over LHC2. Additionally, SSLHC3, which employs multi-sketching, is more efficient than SLHC2 and exhibits better numerical stability. It is also robust as a deterministic algorithm.
Key words and phrases:
QR factorization, Rounding error analysis, LU-CholeskyQR, Numerical linear algebra
2020 Mathematics Subject Classification:
65F35, 15A23, 65F25, 65G50
1Corresponding author. Department of Applied Mathematics, The Hong Kong Polytechnic University, Hung Hom, Hong Kong. Email: 21037226R@connect.polyu.hk. H.Guan’s work is funded by the Hong Kong Research Grants Council under RFS grant RFS2021-5S03 and GRF grant 15302122, as well as by the Hong Kong Polytechnic University grant 4-ZZLS.
2Theory Lab, Huawei Hong Kong Research Center, Sha Tin, Hong Kong. Email: fanyuwei2@huawei.com.
1. Introduction
QR factorization is one of the most important components of numerical linear algebra and is widely used in both academia and industry, see [10, 18, 20, 23, 31, 37] for more details. Among the various algorithms for QR factorization, CholeskyQR has gained popularity in recent years due to its ability to balance efficiency and accuracy. CholeskyQR exclusively utilizes BLAS3 operations and requires only simple reductions in parallel environments, which is a significant advantage compared to other algorithms such as HouseholderQR, CGS (MGS), and TSQR [6, 8, 14, 19, 21, 25].
1.1. Deterministic CholeskyQR-type algorithms
[…]
@@ 1.1. Deterministic CholeskyQR-type algorithms
In practical applications, CholeskyQR is seldom used alone due to its lack of orthogonality. To enhance numerical stability in both orthogonality and residuals, CholeskyQR2 [13, 39] was developed by repeating CholeskyQR twice. This algorithm is outlined in Algorithm 2. To address the issue of ill-conditioned matrices, Shifted CholeskyQR (SCholeskyQR) [12] introduces a shifted term $s$ during the Cholesky factorization step to avoid numerical breakdown, as detailed in Algorithm 3. Researchers have combined CholeskyQR2 with Shifted CholeskyQR to create Shifted CholeskyQR3 (SCholeskyQR3), which ensures orthogonality, as shown in Algorithm 4. In our recent work [11], we define a new $g$-norm of $X$ and improve the original shifted term associated with $\|X\|_{2}$ by using a smaller term based on $\|X\|_{g}$. This modification expands the computational range of $\kappa_{2}(X)$ for Shifted CholeskyQR3. With the definition of the $g$-norm, we also analyze the case when $X$ is sparse [15]. Further analysis of deterministic CholeskyQR-type algorithms can be found in [16, 42, 40].
@@
1.2. LU-CholeskyQR2 and its bottleneck
Among all the deterministic CholeskyQR-type algorithms, LU-CholeskyQR [34] is particularly noteworthy. It has the unique advantage of not imposing a restriction on $\kappa_{2}(X)$, which is crucial for real-world applications in industry. LU-CholeskyQR combines LU factorization with CholeskyQR, as shown in Algorithm 5. Here, $L\in\mathbb{R}^{m\times n}$ is a unit tall-skinny lower-triangular matrix, and $U\in\mathbb{R}^{n\times n}$ is an upper-triangular matrix when $Xin\mathbb{R}^{m\times n}$ is tall-skinny. A CholeskyQR step is then performed to form LU-CholeskyQR2, ensuring orthogonality, as detailed in Algorithm 6.
@@ 1.2. LU-CholeskyQR2 and its bottleneck
In practice, the $L$-factor after LU factorization is often well-conditioned, even when $X$ is ill-conditioned. However, there are many real-world scenarios that can lead to an ill-conditioned $L$, such as certain lower-triangular matrices and diagonally dominant matrices with many negative elements. These situations frequently arise in linear equations, information theory, and systems theory. The primary bottleneck of LU-CholeskyQR2 is the potential for numerical breakdown in these cases, which poses a significant problem that remains to be addressed or improved.
1.3. Randomized algorithms and our contributions
In recent years, randomized numerical linear algebra [18, 23] has become a hot topic worldwide. A new randomized technique called sketching [1] has been widely applied to various problems. The sketching technique primarily aims to reduce computational costs by replacing the original large matrices with alternative matrices of smaller sizes after some preconditioning steps. The sketching matrix is typically chosen to be an $\epsilon$-subspace embedding or a linear map to a lower-dimensional space, preserving the inner products and norms of all vectors within the subspace up to a factor of $1+\epsilon$, where $10\leq\epsilon<$ 1. Several existing methods for sketching matrices include Gaussian Sketch, SRHT, and Count Sketch [5, 22, 26, 33, 4]. Research on randomized CholeskyQR has also emerged. Y. Fan and his collaborators [41] proposed the initial version of randomized CholeskyQR, while O. Balabanov [27] provided a detailed analysis of CholeskyQR-type algorithms with various randomized strategies, including sketching. A recent work by A. J. Higgins and his collaborators [2] introduced a novel method called multi-sketching, which employs two different sketching steps consecutively. This work developed a new algorithm called randomized Householder QR, utilizing HouseholderQR to replace the original generation step of the $R$-factor in CholeskyQR to avoid numerical breakdown. The multi-sketching technique is designed to accelerate the entire algorithm. The corresponding algorithm is outlined in Algorithm 7. Here, $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ and $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ are sketching matrices, with $n\leq s_{2}\leq s_{1}\leq m$. To ensure good orthogonality, a CholeskyQR step is added afterward, resulting in the Rand.Householder-Cholesky algorithm, as shown in Algorithm 8. The authors provided a clear rounding error analysis of the algorithms with two $\epsilon$-subspace embeddings, and numerical experiments demonstrated that their randomized Householder-CholeskyQR is more efficient and numerically stable than the algorithms in [27].
@@ 1.3. Randomized algorithms and our contributions
1.4. Outline of this work
Our work is organized as follows: We present existing results from previous studies in Section 2. In Sections 3 and 4, we introduce our improved deterministic and randomized algorithms, along with their corresponding rounding error analyses concerning orthogonality and residuals. The primary numerical experiments and comparisons are detailed in Section 5. Finally, we conclude with a summary of the results in Section 6.
2. Literature review
@@
2.1. Rounding error analysis
@@
Lemma 2.1 (Weyl’s Theorem [14]).
@@
Lemma 2.2 (Rounding error in matrix multiplications [19]).
For $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times p}$, the error in computing the matrix product $C=AB\in\mathbb{R}^{m\times n}$ in floating-point arithmetic is bounded by
@@
Lemma 2.3 (Rounding error in HouseholderQR [19]).
@@
To learn more about the rounding error analysis of matrix perturbations, readers can refer to [9, 17, 32] for more details.
2.2. Matrix sketching
@@
2.2.1. Single-sketching
In this section, we introduce the concept of sketching with the definition of an $\epsilon$-subspace embedding [5, 26, 29]. Similar results are also mentioned in [2].
Definition 2.4 ($\epsilon$-subspace embedding).
@@ 2.2.1. Single-sketching
(2.1) $|\langle x,y\rangle-\langle\Omega x,\Omega y\rangle|\leq\epsilon\|x\|_{2}\|y\|_{2}.$
@@
In the following, we list some properties based on Definition 2.4.
Lemma 2.5 (Single-sketching of the vector).
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, then for any $x\in\mathcal{K}$, we have
(2.2) $\sqrt{1-\epsilon}\cdot\|x\|_{2}\leq\|\Omega x\|_{2}\leq\sqrt{1+\epsilon}\cdot\|x\|_{2}.$
Lemma 2.5 focuses on the sketching of vectors. Below, we list the properties of the matrix after sketching.
Lemma 2.6 (Single-sketching of the matrix).
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, if $T$ is a matrix whose columns form a basis of $\mathcal{K}$, we have
(2.3) $\sqrt{1-\epsilon}\cdot\|T\|_{2}\leq\|\Omega T\|_{2}\leq\sqrt{1+\epsilon}\cdot\|T\|_{2},$
(2.4) $\sqrt{1-\epsilon}\cdot\|T\|_{F}\leq\|\Omega T\|_{F}\leq\sqrt{1+\epsilon}\cdot\|T\|_{F},$
(2.5) $\sqrt{1-\epsilon}\cdot\sigma_{min}(T)\leq\sigma_{min}(\Omega T)\leq\|\Omega T\|_{2}\leq\sqrt{1+\epsilon}\cdot\|T\|_{2}.$
@@
In Lemma 2.6, (2.3) and (2.5) play a crucial role in our analysis. (2.4) can be easily derived from (2.3). Below, we list some existing results related to condition numbers.
Lemma 2.7 (Condition number of the single-sketching).
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-subspace embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, if $T$ is a matrix whose columns form a basis of $\mathcal{K}$, we have
(2.6) $\frac{\sigma_{min}(\Omega T)}{\sqrt{1+\epsilon}}\leq\sigma_{min}(T)\leq\|T\|_{2}\leq\frac{\|\Omega T\|_{2}}{\sqrt{1-\epsilon}}.$
That is:
(2.8) $\kappa_{2}(T)\leq\sqrt{\frac{1-\epsilon}{1+\epsilon}}\cdot\kappa_{2}(\Omega T).$
@@
The results mentioned above pertain to $\epsilon$-subspace embeddings, which require knowledge of the subspace $\mathcal{K}$. In contrast, a concept known as the $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding [5] does not require this knowledge.
Definition 2.8 ($(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding).
@@
Typical $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embeddings include both dense and sparse cases, which correspond to different sketching methods. A sketching matrix is defined as $\Omega=\frac{1}{s}G$, where $G\in\mathbb{R}^{s\times m}$ is a Gaussian matrix. The sketching size $s$ [29] satisfies
(2.9) $s=\eta\cdot\frac{logn\cdot log{\frac{1}{p}}}{\epsilon^{2}}.$
@@
Regarding $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embeddings, researchers often use CountSketch [24]. If $\Omega\in\mathbb{R}^{s\times m}$ is a sketching matrix for CountSketch, then, according to [7, 38], $s$ needs to satisfy
(2.10) $s\geq\frac{n^{2}+n}{\epsilon^{2}p}.$
2.2.2. Multi-sketching
In [2], a technique called multi-sketching is introduced, which involves applying two steps of sketching consecutively. Below, we list some properties of the multi-sketching technique.
Lemma 2.9 (Multi-sketching of the vector).
If $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ is a $(\epsilon_{2},p_{2},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$ and they are generated independently, then for any $n$-dimensional subspace $\mathcal{K}\subset\mathbb{R}^{m}$ and any $x\in\mathcal{K}$, we have
(2.11) $\sqrt{1-\epsilon_{s}}\cdot\|x\|_{2}\leq\|\Omega_{2}\Omega_{1}x\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|x\|_{2}$